# The Enduring Relevance of Semiempirical Quantum Mechanics

### v2

@@ 1 Introduction
While SQM method development has continued to the present day, it has been overshadowed in accuracy by ab initio QM methods and in efficiency by molecular mechanics (MM) methods. The growing abundance of computational resources has enabled the widespread use of QM methods, and the efficiency needs of large-scale atomistic simulations have only been satisfied by MM methods. In general, a polarization of method development in atomistic simulation between accurate QM methods and fast MM methods has been driven by the pursuit of specific challenging applications that demand either accuracy or efficiency beyond what computational methods have been able to deliver. In quantum chemistry, this polarization 13, 14, 15 was driven by the pursuit of a fully predictive QM-based understanding of small molecules and a more descriptive atomistic understanding of biological processes by embedding small QM regions in large MM simulations and coupling them with QM/MM methods 16. Modern QM method development has focused on density functional theory 17 (DFT) to provide a baseline level of accuracy that might then be refined by more expensive quantum many-body methods such as coupled-cluster theory 18 or many-body Green’s functions 19. Although the AM1 model paper is among the most cited scientific papers, several important DFT methodology papers collectively have an order of magnitude more citations 20. Practical DFT calculations retain several important SQM influences, mainly the use of semiempirical density functionals such as B3LYP 21, the evolution of EPM models into the plane-wave pseudopotential formalism widely used in physics 22, and the generation of system-specific tight-binding models as a common post-processing step of plane-wave pseudopotential calculations 23. Despite these influences, a large gap in computational cost continues to separate QM from SQM methods, and a similarly large gap separates SQM from MM methods. For the elementary task of atomic force evaluation, the relative computational cost separating QM, SQM, and MM calculations is typically larger than one thousand 24. Thus, I use the QM/SQM/MM categorization in this paper to denote regimes of computational cost necessitated by the level of theory.
@@ 2 Models from components
Historically, the most popular SQM software has been MOPAC 30, which implemented the MNDO family 31 of SQM models including AM1 7. The popularity of DFT in the 1990’s inspired the density functional tight binding (DFTB) method 32 and eventually the DFTB+ software 33. More recently, the Grimme group at the University of Bonn developed the GFN family of SQM models and its software implementation in xTB 34. These programs have been the primary platforms for the development and use of their respective SQM models, and I visualize their citation data in Fig. 2. Other impactful examples of SQM models and software are Zerner’s INDO model and ZINDO program 35, the Naval Research Laboratory (NRL) tight-binding model 36, and the NEMO tight-binding model and program 37.
While not a strict separation, SQM software is typically intended for specific computational tasks with a limited cost while QM software is typically intended to exhaust available computing resources to achieve as much certainty as possible, as exemplified by Pople’s diagrams 38 and Perdew’s ladder 39. For example, MOPAC predicts the heat of formation and equilibrium geometry of molecules, xTB is heavily used with the CREST software 40 for conformational sampling of molecules, ZINDO predicts the electronic spectra of molecules, and NEMO is used for semiconductor device modeling, while DFTB+ and the NRL tight-binding model are more generically used as low-cost approximations of DFT calculations. Both QM and SQM software have numerical parameters that can be tuned, but only QM software effectively enables a systematic reduction of all sources of error if sufficient computational resources are available. While the computational cost of individual MM calculations is much lower than QM and SQM calculations, MM software is typically used for tasks that require a very large number of calculations to control bias and reduce uncertainty of statistical estimates, and available computing resources are often exhausted to that end. Similarly, SQM software provides the most cost-effective access to large quantities of electronic properties that are not provided by MM calculations, for ensemble averaging as in CREST or for high-throughput virtual screening of molecules and materials 41.
@@
To the degree that QM methods and software are modular, they can be considered as components for the construction of new SQM models. In this way, some low-cost use of QM software could avoid the human cost of calibration for common atomistic simulation tasks. Indeed, this has been a popular activity in modern QM method development with particularly notable results from Grimme such as semiempirical density functionals 42, the D models 43 for dispersion corrections to popular density functionals, and the -3c models 44 that correct both basis-set incompleteness and correlation errors. However, in using components from modern QM software, these SQM models do not have access to the low-cost components that enable the efficiency of dedicated SQM software. The efficiency of AO-based SQM software depends on the neglect of diatomic differential overlap (NDDO) approximation, which is an approximate factorization of four-center Coulomb integrals that was introduced during the development of the CNDO model 45 but still has no modern QM counterpart. Similarly, the efficiency of EPM models depends on very soft pseudopotentials that enable a very small plane-wave basis with a very low energy cutoff, which was lost in the transition towards harder, more transferable pseudopotential forms in modern plane-wave pseudopotential QM methods.
@@
The ability for QM software to be reshaped into new SQM models depends on how effectively it has been decomposed into reliable and reconfigurable components, which is a steadily growing emphasis of atomistic simulation software development 46. Such components are a tightly coupled stack of interfaces (e.g., user interface, application programming interface, input and output data formats) on top of software, algorithms, and methods, resting on foundations of theoretical formalism. A substantial amount of recent activity has focused on the interface and software layers of this component stack. First, there are a few consolidated efforts to decompose QM software into more independent components such as the PySCF program 47 implemented in the highly modular Python programming language and the ESL library bundle 48 that encapsulates most of the functionality of a plane-wave pseudopotential QM program. Next, there are a variety of interfacing layers to enable existing software to function as interoperable components such as the high-level SEAMM environment 49 that enables the construction of multi-program workflows as graphical flowcharts and the low-level MDI library 50 for run-time inter-program communication. Data is the most ubiquitous interface between programs and users, and there is active movement towards findable, accessible, interoperable, and reusable (FAIR) data standards throughout atomistic simulation 51 and important specific instantiations like the Basis Set Exchange 52 for Gaussian basis set data and the QCSchema 53 that enables interoperable use of quantum chemistry software in databases and workflow systems. A growing number of critical components of QM software now have software libraries, including Gaussian AO integrals 54, DFT exchange-correlation functionals 55, Hamiltonian matrix solvers 56, molecular geometry optimizers 57, and self-consistent field (SCF) solvers 58.
@@
While interfaces and software are major parts of developing atomistic simulation components, refinement of the underlying algorithms, methods, and even formalism may also be important in some cases. Low-level problems in a component stack can be mitigated by interface and software design such as error handling, but they cannot be fixed without resolving the problem at its lowest level. Without some active form of error correction, reliable QM software needs reliable components. By far, the least reliable part of any QM software is the SCF cycle, because development has focused on algorithms such as direct inversion in the iterative subspace (DIIS) 59 that reduce costs for typical problem instances in chemistry rather than reduce the cost or prevalence of difficult problem instances that suffer from convergence pathologies. SCF solvers have been attracting increasing attention from applied mathematicians 60, who have demonstrated and proven the increased reliability of direct minimization SCF solvers over mixing algorithms like DIIS. However, SCF solver algorithms based on direct minimization tend to be slower for typical problem instances, and it is not yet clear if this is a temporary or fundamental limitation. Reliable SCF solvers may also need active electronic orbitals with fractional occupations in some problem instances 61, which is not directly compatible with the strict partition of electronic orbitals into occupied and virtual in many quantum chemistry programs. Ideally, developer culture in atomistic simulation should align software and method development concerns in the transition towards more component-centric development, with software support activity focused on identifying practical priorities for future method development and software maintenance activity focused on implementing new methods and integrating new knowledge as they emerge.
@@ 2.1 SQM model deconstruction
While the historical development of SQM models has been monolithic, their many design choices can be tabulated and compared 29. In addition to differing in how they are fit and what they are fit to, these design choices primarily cluster into three components: a model Hamiltonian, a total energy model, and a solver algorithm. Improved modularity of these components are important to understanding their individual costs and errors. SQM models are intentionally cost-limited approximations, which makes it particularly important to understand how costs and errors are budgeted between their different components. A lack of such understanding can imbalance costs and errors, reducing the overall effectiveness of an SQM model. Historical examples of this include the limited effectiveness of combining second-order many-body perturbation theory with MNDO-family models 62, and the persistent failure of MNDO-family models to describe the electrostatics of hydrogen bonds 63 without explicit classical corrections 64. Building SQM models from more modular components also enables a separation of concerns between improving components and designing new SQM models that use them.
@@
In isolation, the fitting of a model many-electron Hamiltonian in SQM has the same goal as more formal QM methods for transforming ab initio Hamiltonians into simpler model forms 65. The main difference is that SQM model Hamiltonians are parameterized to fit reference data directly rather than carrying out an explicit many-body operator transformation. These research activities have a lot to learn from each other, as SQM has been successfully deploying specific parameterized families of model Hamiltonians for decades while many-body transformations offer deeper insight into the effectiveness of different Hamiltonian model forms. More specifically, efforts to increase the accuracy of SQM models are now going beyond a minimal AO basis and following the popular quantum chemistry prescription of including split-valence and polarization functions 66. However, increasing the number of AOs per atom tends to increase the condition number of the AO overlap matrix, which may have more severe consequences for SQM models than ab initio QM calculations. A possible alternative might be to revisit EPM model forms 10 to construct SQM model Hamiltonians with both a minimal AO basis and a set of smooth, uniform basis functions with a similar number of functions per atom as a larger AO basis. The resulting hybrid SQM model Hamiltonians would have new classes of matrix elements to parameterize such as matrix elements between AOs and low-degree polynomials.
@@
The choice of total energy components for SQM models is slowly expanding with increasing progress in QM method development. The MNDO-family models in MOPAC 31 are based on Hartree-Fock calculations with total energies fit to reproduce experimental heats of formation because that was the efficient level of theory and accessible reference data available in its time. The later DFTB models 32 used expansions around an atomic limit to construct total energy models that approximate ab initio DFT calculations, which then served to supplement experimental reference data. As independent components, a model Hamiltonian and total energy model should both be as accurate as possible in isolation, although it may still be beneficial to reparameterize model Hamiltonians to compensate for errors in the accompanying total energy model. Fast algorithms can enable more accurate total energy models in cost-constrained SQM models, and recent innovations include cubic-scaling algorithms for the random-phase approximation (RPA) and second-order Møller-Plesset perturbation theory 67. While SQM models tend to focus on ground-state properties, there are also advances in total energy models for excited states such as the use of quasiparticle methods with tight-binding models 68. Because of the broad interest in the quantum many-body problem, new total energy components may come from research areas outside of atomistic simulation, such as quantum information theory, nuclear physics, or high-energy particle physics.
@@
To achieve a low computational cost, SQM models may need to use solver algorithms that either introduce additional approximations and errors or reduce the domain of applicability, or else wait for future methodological advances. There was a substantial amount of research in the 1990s on linear-scaling QM solvers 69, and SQM models were the first beneficiaries of this research because they could effectively use these solvers for smaller system sizes than ab initio QM applications. The MOZYME solver in MOPAC 70 was one of the earliest linear-scaling solvers with a practical deployment because its application domain is restricted to closed-shell systems with identifiable formal atomic charges and bond orders. More generally applicable linear-scaling solvers are now available to SQM models 71, but they introduce errors that must be budgeted into the overall SQM model. There is a broad range of linear-scaling solvers available to SQM models, with the most efficient and approximate solvers having natural connections to bond-order potentials used in MM methods 72. While linear-scaling solvers with both high accuracy and general applicability do not yet exist, progress is slowly being made in numerical linear algebra research towards that end. For example, the pole expansion and selected inversion (PEXSI) method 73 is a quadratic-scaling QM solver that utilizes modern advances in sparse matrix factorization and is not restricted to closed-shell systems. Applied mathematicians have known for over three decades how to compress the free electron propagator matrix with fast multipole methods 74, but those fundamental advances have still not been generalized to electrons in inhomogeneous molecules and materials.
@@ 2.2 Statistical model selection
Once its components have been selected, an SQM model is constructed by fitting its parameters to reference data. Typically, the fitting process is carried out as a nonlinear least-squares optimization 75, such as is implemented in the PARAM fitting program that accompanies MOPAC. However, this process does not provide any operational meaning to a model’s goodness of fit, nor does it provide feedback on the selection of model components or protection against overfitting. All of these benefits can be obtained by applying modern statistics and its formalism for model selection to the construction of SQM models 76. This is a maximalist approach that comes with many technical requirements, but it provides a precise statistical meaning to constructing the SQM model that is most likely to succeed at a simulation task, provided that a precise statistical description for a desired distribution of tasks is constructed first.
@@
In the maximalist interpretation, a set of reference data corresponds to independent samples from a target distribution of atomistic simulation tasks. For example, the experimental heat-of-formation data used to parameterize the PM models in MOPAC are direct samples from the distribution of molecules that are of experimental interest. QM method development has been steadily aggregating its own large reference data sets such as MGCDB84 77, but it is less clear how they relate to task distributions of practical applications. Reference data serves as more of a satisfiability test in QM method development because such methods are expected to have a high degree of transferability, therefore they can be expected to be accurate in general if they are accurate in a sufficient number of arbitrary examples. The effectiveness of SQM models depends on the similarity of testing and training sets, such as in the case of GFN models performing better than PM models for conformer energy prediction 78 because GFN models are more directly trained on intermolecular interaction energies. With the ongoing exponential growth in the use of atomistic simulation methods and software, it is becoming increasingly important to understand this use better and characterize it with a set of task distributions corresponding to the most prevalent patterns of use. The work of constructing such task distributions is distinct from QM method development or SQM model fitting, but it can help to target those activities more effectively at real demand. Even ab initio QM method development is dependent on task distributions, with molecular applications dominated by the Gaussian AO formalism and materials applications dominated by the plan-wave pseudopotential formalism.
@@
An additional benefit of having well-developed task distributions and success metrics is that they can be used to score and rank different methods and software. Rankings can offset the confusion and decision paralysis of a large number of choices, such as with Gaussian AO basis sets or DFT exchange-correlation functionals. Leaderboard websites for ranking atomistic simulation methods and software are already beginning to emerge 79. If rankings include the average computational cost alongside the success metric, then a Pareto front of cost versus accuracy can be constructed for the available software. Significant gaps in these Pareto fronts can then identify deficiencies in atomistic simulation capabilities and useful targets for new method and software development. Such a reduction of diverse methods and software to cost-accuracy curves for a set of task distributions is a step towards the commoditization of atomistic simulation.

### v3

@@ 1 Introduction
While SQM method development has continued to the present day, it has been overshadowed in accuracy by ab initio QM methods and in efficiency by molecular mechanics (MM) methods. The growing abundance of computational resources has enabled the widespread use of QM methods, and the efficiency needs of large-scale atomistic simulations have only been satisfied by MM methods. In general, a polarization of method development in atomistic simulation between accurate QM methods and fast MM methods has been driven by the pursuit of specific challenging applications that demand either accuracy or efficiency beyond what computational methods have been able to deliver. In quantum chemistry, this polarization 13, 14, 15 was driven by the pursuit of a fully predictive QM-based understanding of small molecules and a more descriptive atomistic understanding of biological processes by embedding small QM regions in large MM simulations and coupling them with QM/MM methods 16. Modern QM method development has focused on density functional theory 17 (DFT) to provide a baseline level of accuracy that might then be refined by more expensive quantum many-body methods such as coupled-cluster theory 18 or many-body Green’s functions 19. Although the AM1 model paper is among the most cited scientific papers, several important DFT methodology papers collectively have an order of magnitude more citations 20. Practical DFT calculations retain several important SQM influences: the use of semiempirical density functionals such as B3LYP 21, the evolution of EPM models into the plane-wave pseudopotential formalism widely used in physics 22, and the generation of system-specific tight-binding models as a common post-processing step of plane-wave pseudopotential calculations 23. Despite these influences, a large gap in computational cost continues to separate QM from SQM methods, and a similarly large gap separates SQM from MM methods. For the elementary task of atomic force evaluation, the relative computational cost separating QM, SQM, and MM calculations is typically larger than one thousand 24. Thus, I use the QM/SQM/MM categorization in this paper to denote regimes of computational cost necessitated by the level of theory.
@@ 2 Models from components
Historically, the most popular SQM software has been MOPAC 30, which implemented the MNDO family 6 of SQM models including AM1 7. The popularity of DFT in the 1990’s inspired the density functional tight binding (DFTB) method 31 and eventually the DFTB+ software 32. More recently, the Grimme group at the University of Bonn developed the GFN family of SQM models and its software implementation in xTB 33. These programs have been the primary platforms for the development and use of their respective SQM models, and I visualize their citation data in Fig. 2. Other impactful examples of SQM models and software are Zerner’s INDO model and ZINDO program 34, the Naval Research Laboratory (NRL) tight-binding model 35, and the NEMO tight-binding model and program 36.
While not a strict separation, SQM software is typically intended for specific computational tasks with a limited cost while QM software is typically intended to exhaust available computing resources to achieve as much certainty as possible, as exemplified by Pople’s diagrams 37 and Perdew’s ladder 38. For example, MOPAC predicts the heat of formation and equilibrium geometry of molecules, xTB is heavily used with the CREST software 39 for conformational sampling of molecules, ZINDO predicts the electronic spectra of molecules, and NEMO is used for semiconductor device modeling, while DFTB+ and the NRL tight-binding model are more generically used as low-cost approximations of DFT calculations. Both QM and SQM software have numerical parameters that can be tuned, but only QM software effectively enables a systematic reduction of all sources of error if sufficient computational resources are available. While the computational cost of individual MM calculations is much lower than QM and SQM calculations, MM software is typically used for tasks that require a very large number of calculations to control bias and reduce uncertainty of statistical estimates, and available computing resources are often exhausted to that end. Similarly, SQM software provides the most cost-effective access to large quantities of electronic properties that are not provided by MM calculations, for ensemble averaging as in CREST or for high-throughput virtual screening of molecules and materials 40.
@@
To the degree that QM methods and software are modular, they can be considered as components for the construction of new SQM models. In this way, some low-cost use of QM software could avoid the human cost of calibration for common atomistic simulation tasks. Indeed, this has been a popular activity in modern QM method development with particularly notable results from Grimme such as semiempirical density functionals 41, the D models 42 for dispersion corrections to popular density functionals, and the -3c models 43 that correct both basis-set incompleteness and correlation errors. However, in using components from modern QM software, these SQM models do not have access to the low-cost components that enable the efficiency of dedicated SQM software. The efficiency of AO-based SQM software depends on the neglect of diatomic differential overlap (NDDO) approximation, which is an approximate factorization of four-center Coulomb integrals that was introduced during the development of the CNDO model 44 but still has no modern QM counterpart. Similarly, the efficiency of EPM models depends on very soft pseudopotentials that enable a very small plane-wave basis with a very low energy cutoff, which was lost in the transition towards harder, more transferable pseudopotential forms in modern plane-wave pseudopotential QM methods.
@@
The ability for QM software to be reshaped into new SQM models depends on how effectively it has been decomposed into reliable and reconfigurable components, which is a steadily growing emphasis of atomistic simulation software development 45. Such components are a tightly coupled stack of interfaces (e.g., user interface, application programming interface, input and output data formats) on top of software, algorithms, and methods, resting on foundations of theoretical formalism. A substantial amount of recent activity has focused on the interface and software layers of this component stack. First, there are a few consolidated efforts to decompose QM software into more independent components such as the PySCF program 46 implemented in the highly modular Python programming language and the ESL library bundle 47 that encapsulates most of the functionality of a plane-wave pseudopotential QM program. Next, there are a variety of interfacing layers to enable existing software to function as interoperable components such as the high-level SEAMM environment 48 that enables the construction of multi-program workflows as graphical flowcharts and the low-level MDI library 49 for run-time inter-program communication. Data is the most ubiquitous interface between programs and users, and there is active movement towards findable, accessible, interoperable, and reusable (FAIR) data standards throughout atomistic simulation 50 and important specific instantiations like the Basis Set Exchange 51 for Gaussian basis set data and the QCSchema 52 that enables interoperable use of quantum chemistry software in databases and workflow systems. A growing number of critical components of QM software now have software libraries, including Gaussian AO integrals 53, DFT exchange-correlation functionals 54, Hamiltonian matrix solvers 55, molecular geometry optimizers 56, and self-consistent field (SCF) solvers 57.
@@
While interfaces and software are major parts of developing atomistic simulation components, refinement of the underlying algorithms, methods, and even formalism may also be important in some cases. Low-level problems in a component stack can be mitigated by interface and software design such as error handling, but they cannot be fixed without resolving the problem at its lowest level. Without some active form of error correction, reliable QM software needs reliable components. By far, the least reliable part of any QM software is the SCF cycle, because development has focused on algorithms such as direct inversion in the iterative subspace (DIIS) 58 that reduce costs for typical problem instances in chemistry rather than reduce the cost or prevalence of difficult problem instances that suffer from convergence pathologies. SCF solvers have been attracting increasing attention from applied mathematicians 59, who have demonstrated and proven the increased reliability of direct minimization SCF solvers over mixing algorithms like DIIS. However, SCF solver algorithms based on direct minimization tend to be slower for typical problem instances, and it is not yet clear if this is a temporary or fundamental limitation. Reliable SCF solvers may also need active electronic orbitals with fractional occupations in some problem instances 60, which is not directly compatible with the strict partition of electronic orbitals into occupied and virtual in many quantum chemistry programs. Ideally, developer culture in atomistic simulation should align software and method development concerns in the transition towards more component-centric development, with software support activity focused on identifying practical priorities for future method development and software maintenance activity focused on implementing new methods and integrating new knowledge as they emerge.
@@ 2.1 SQM model deconstruction
While the historical development of SQM models has been monolithic, their many design choices can be tabulated and compared 29. In addition to differing in how they are fit and what they are fit to, these design choices primarily cluster into three components: a model Hamiltonian, a total energy model, and a solver algorithm. Improved modularity of these components is important to understanding their individual costs and errors. SQM models are intentionally cost-limited approximations, which makes it particularly important to understand how costs and errors are budgeted between their different components. A lack of such understanding can imbalance costs and errors, reducing the overall effectiveness of an SQM model. Historical examples of this include the limited effectiveness of combining second-order many-body perturbation theory with MNDO-family models 61, and the persistent failure of MNDO-family models to describe the electrostatics of hydrogen bonds 62 without explicit classical corrections 63. Building SQM models from more modular components also enables a separation of concerns between improving components and designing new SQM models that use them.
@@
In isolation, the fitting of a model many-electron Hamiltonian in SQM has the same goal as more formal QM methods for transforming ab initio Hamiltonians into simpler model forms 64. The main difference is that SQM model Hamiltonians are parameterized to fit reference data directly rather than carrying out an explicit many-body operator transformation. These research activities have a lot to learn from each other, as SQM has been successfully deploying specific parameterized families of model Hamiltonians for decades while many-body transformations offer deeper insight into the effectiveness of different Hamiltonian model forms. More specifically, efforts to increase the accuracy of SQM models are now going beyond a minimal AO basis and following the popular quantum chemistry prescription of including split-valence and polarization functions 65. However, increasing the number of AOs per atom tends to increase the condition number of the AO overlap matrix, which may have more severe consequences for SQM models than ab initio QM calculations. A possible alternative might be to revisit EPM model forms 10 to construct SQM model Hamiltonians with both a minimal AO basis and a set of smooth, uniform basis functions with a similar number of functions per atom as a larger AO basis. The resulting hybrid SQM model Hamiltonians would have new classes of matrix elements to parameterize such as matrix elements between AOs and low-degree polynomials.
@@
The choice of total energy components for SQM models is slowly expanding with increasing progress in QM method development. The MNDO-family models in MOPAC 6 are based on Hartree-Fock calculations with total energies fit to reproduce experimental heats of formation because that was the efficient level of theory and accessible reference data available in its time. The later DFTB models 31 used expansions around an atomic limit to construct total energy models that approximate ab initio DFT calculations, which then served to supplement experimental reference data. As independent components, a model Hamiltonian and total energy model should both be as accurate as possible in isolation, although it may still be beneficial to reparameterize model Hamiltonians to compensate for errors in the accompanying total energy model. Fast algorithms can enable more accurate total energy models in cost-constrained SQM models, and recent innovations include cubic-scaling algorithms for the random-phase approximation (RPA) and second-order Møller-Plesset perturbation theory 66. While SQM models tend to focus on ground-state properties, there are also advances in total energy models for excited states such as the use of quasiparticle methods with tight-binding models 67. Because of the broad interest in the quantum many-body problem, new total energy components may come from research areas outside of atomistic simulation, such as quantum information theory, nuclear physics, or high-energy particle physics.
@@
To achieve a low computational cost, SQM models may need to use solver algorithms that either introduce additional approximations and errors or reduce the domain of applicability, or else wait for future methodological advances. There was a substantial amount of research in the 1990s on linear-scaling QM solvers 68, and SQM models were the first beneficiaries of this research because they could effectively use these solvers for smaller system sizes than ab initio QM applications. The MOZYME solver in MOPAC 69 was one of the earliest linear-scaling solvers with a practical deployment because its application domain is restricted to closed-shell systems with identifiable formal atomic charges and bond orders. More generally applicable linear-scaling solvers are now available to SQM models 24, but they introduce errors that must be budgeted into the overall SQM model. There is a broad range of linear-scaling solvers available to SQM models, with the most efficient and approximate solvers having natural connections to bond-order potentials used in MM methods 70. While linear-scaling solvers with both high accuracy and general applicability do not yet exist, progress is slowly being made in numerical linear algebra research towards that end. For example, the pole expansion and selected inversion (PEXSI) method 71 is a quadratic-scaling QM solver that utilizes modern advances in sparse matrix factorization and is not restricted to closed-shell systems. Applied mathematicians have known for over three decades how to compress the free electron propagator matrix with fast multipole methods 72, but those fundamental advances have still not been generalized to electrons in inhomogeneous molecules and materials.
@@ 2.2 Statistical model selection
Once its components have been selected, an SQM model is constructed by fitting its parameters to reference data. Typically, the fitting process is carried out as a nonlinear least-squares optimization 73, such as is implemented in the PARAM fitting program that accompanies MOPAC. However, this process does not provide any operational meaning to a model’s goodness of fit, nor does it provide feedback on the selection of model components or protection against overfitting. All of these benefits can be obtained by applying modern statistics and its formalism for model selection to the construction of SQM models 74. This is a maximalist approach that comes with many technical requirements, but it provides a precise statistical meaning to constructing the SQM model that is most likely to succeed at a simulation task, provided that a precise statistical description for a desired distribution of tasks is constructed first.
@@
In the maximalist interpretation, a set of reference data corresponds to independent samples from a target distribution of atomistic simulation tasks. For example, the experimental heat-of-formation data used to parameterize the PM models in MOPAC are direct samples from the distribution of molecules that are of experimental interest. QM method development has been steadily aggregating its own large reference data sets such as MGCDB84 75 but it is less clear how they relate to task distributions of practical applications. Reference data serves as more of a satisfiability test in QM method development because such methods are expected to have a high degree of transferability, therefore they can be expected to be accurate in general if they are accurate in a sufficient number of arbitrary examples. The effectiveness of SQM models depends on the similarity of testing and training sets, such as in the case of GFN models performing better than PM models for conformer energy prediction 76 because GFN models are more directly trained on intermolecular interaction energies. With the ongoing exponential growth in the use of atomistic simulation methods and software, it is becoming increasingly important to understand this use better and characterize it with a set of task distributions corresponding to the most prevalent patterns of use. The work of constructing such task distributions is distinct from QM method development or SQM model fitting, but it can help to target those activities more effectively at real demand. Even ab initio QM method development is dependent on task distributions, with molecular applications dominated by the Gaussian AO formalism and materials applications dominated by the plane-wave pseudopotential formalism.
@@
An additional benefit of having well-developed task distributions and success metrics is that they can be used to score and rank different methods and software. Rankings can offset the confusion and decision paralysis of a large number of choices, such as choosing Gaussian AO basis sets or DFT exchange-correlation functionals. Leaderboard websites for ranking atomistic simulation methods and software are already beginning to emerge 77. If rankings include the average computational cost alongside the success metric, then a Pareto front of cost versus accuracy can be constructed for the available software. Significant gaps in these Pareto fronts can then identify deficiencies in atomistic simulation capabilities and useful targets for new method and software development. Such a reduction of diverse methods and software to cost-accuracy curves for a set of task distributions is a step towards the commoditization of atomistic simulation.